Reacted with 20.0g oxygen and produces 38170, j of
Answer #1

5)

a) Given,

Heat added = 2000 kJ

Mass of MgCO3 = 1800 g

We know that Molar Mass of MgCO3 = 84.3 g /mol

=> Moles of MgCO3 = 1800 / 84.3 = 21.35 moles

MgCO3 ----> MgO + CO2 ....... delta H = 117.3 kJ

=> To break 1 mole of MgCO3 completely into MgO and CO2, 117.3 kJ heat is required

=> Number of moles of MgCO3 that can be decomposed by 2000 kJ of heat = 2000 / 117.3 = 17.05 moles

=> Moles of MgCO3 not dissociated = 21.35 - 17.05 = 4.3 moles

According to stoichiometry of the reaction 1 mole of MgCO3 decomposes to produce 1 mole of MgO and 1 mole of CO2

=> Moles of CO2 produced = 17.05

b)

Moles of MgCO3 remaining = 4.3

=> Mass of MgCO3 = 4.3 x 84.3 = 362.5 g

Moles of MgO produced = 17.05

=> Mass of MgO = 17.05 x 40.3 = 687.12 g

=> Total mass of MgO + MgCO3 = 362.5 + 687.12 = 1049.62 g